Previous

Calendar

October

2023

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

1

JD274

2

JD275
Yom Kippur

3

JD276
light snow (night of 10/2)

4

JD277
AmeriFlux Annual Meeting (Harvard Forest)

5

JD278
AmeriFlux Annual Meeting (Harvard Forest)

6

JD279
Site Visit (SpB): Details , Photos (57)

Span cylinder at 1000 psi, N2 at 800 psi

removed the PolaRx5 GPS "A" (sn 3062189) and antenna from the top of the tower

downloaded PolaRx5 GPS "B" (sn 3227662) data files (in forest); unplugged and removed gpsB antenna

at 11:05 MST, swapped SD card from GPSCAM

Powered on MRS/Noah loggers at N. Canopy Tower (for snowdepths)

downloaded the AP200 data

used the LI-850 software to confirm that the span/zero coeffients were unchanged

at around mid-day the 21.5m LI-7200 inlet flow was clogged (due to humidity?)

Enterprise SUV rental

In-cloud/fog all day; temp is just-above freezing

7

JD280

8

JD281

9

JD282
Indigenous Peoples Day

10

JD283
at around mid-day the 21.5m LI-7200 inlet flow became un-clogged (took 4 days!)

11

JD284

12

JD285
First signigifcant snow of season (approx 3-4 in.)

13

JD286

14

JD287

15

JD288

16

JD289
Site Visit (SpB): Details , Photos (38)

Span cylinder at 1000 psi, N2 at 750 psi

moved 3 cylinders from MRS snowshed to C-1; cc476841 (n2), cc301531 (n2), cc153288 (co2)

moved small LBL span cylinder (ru000402) from C-1 to tower (to use with AP200)

from 10:29-10:35 MST swapped the 21.5m LI-7200 inlet (not completely blocked?)

removed the PolaRx5 GPS "B" (sn 3227662) "forest" receiver and antenna tripod (to return to UNAVCO)

from 12:10-12:20 MST, tried to re-level the CNR4 radiometer

at 13:00 MST, added E-type thermocouple to to path of 2.5m CSAT3 (now using uc_10.csi)

started ops108 (added Ttc.2.5m) to DSM

downloaded the AP200 data; running new program (AP200_spb09.CR1X); added 2 temperature sensors (SE 1/2), connected blue filter to ZERO-air inlet

Enterprise SUV rental

Clear skies, sunny, warm, calm.

17

JD290

18

JD291

19

JD292

20

JD293
in PM, lost connection to the tower (this week urquell data has been missing many samples)

[fiber-optic line to C-1 was damaged]

21

JD294

22

JD295

23

JD296
Site Visit (SpB): Details , Photos (33)

Span cylinder at 950 psi, N2 at 700 psi

moved a bunch of stuff from C-1 trailer to campus

removed the last of the UNAVCO GPS stuff

turned on the MRS logger (snowprobes, T from upper levels in probeA (tree) seem way off!)

for AP200: added 4 more temperature sensors setup as a comparison (now have SE 1/2/3/4/7/8), connected N2 to ZERO-air inlet; moved blue filter to span inlet; connected the heating cable to the upper-four inlets

network connection still down (and was down all weekend); at MRS, the network was ok.

Enterprise SUV rental

Warm, calm.

24

JD297

25

JD298
connection to tower still down (sent OIT email)

26

JD299

27

JD300
fiber-optic line to C-1 is repaired (Kris/MRS arranged fix)

at around 15 MST, the network connection is back up!

28

JD301
(Full "Hunters" Moon)

snow

29

JD302
Signigifcant snow (approx 7-8 in.)

30

JD303
isffa battery voltage starts decreasing

Cold

31

JD304
Halloween

1

JD305
isffa battery voltage cycling (11.5 - 13 V)

2

JD306

3

JD307
Site Visit (SpB, HR): Details , Photos (19)

Span cylinder at 800 psi, N2 at 500 psi

moved n2 cylinder cc236580 from C-1 trailer to tower (outside shed)

moved n2 cylinder cc254426 from tower to C-1 trailer

connected regulator to small span cylinder ru000402 (left at 1700 psi)

brought up new zip ties, bungee cords

at ? MST, swapped SD card from GPSCAM (even though gps is no longer there!)

for AP200: added final 2 temperature sensors (now have SE 1-8), connected the heating cable to the lower four inlets

Enterprise rental

Breezy and cool

4

JD308
the LI-6262 cals seem slighly modified (perhaps due to AP200?)